# DisCo: Disentangled Control for Referring Human Dance Generation in Real World

### v1

@@ 3.3 Human Attribute Pre-training
In utilizing the disentangled control architecture for DisCo, although it shows promises in pose control and background reconstruction, we find it remains challenging to have faithful generations with unseen human subject foregrounds, demonstrating poor generalizability. The crux of this matter lies in the current training pipeline. It relies on high-quality human videos to provide training pairs of human images, with the same human foreground and background appearance, but different poses. Yet, we observe that current training datasets used for human dance generation confront a dilemma of “mutual exclusivity” — they cannot ensure both diversity in human attributes (such as identity, clothing, makeup, hairstyle, etc.) and the complicated poses due to the prohibitive costs of collecting and filtering human videos. As an alternative, human images, which are widely available over the internet, contain diverse human subject foregrounds and backgrounds, despite of the missing paired images with pose alterations.
This motivates us to propose a pre-training task with DisCo, human attribute pre-training, to improve the generalizability and the faithfulness in generation when encountering unseen human subjects. Rather than directly retrieving and constructing the high-quality human dance video dataset, we explore a much more efficient alternative approach, namely Human Attribute Pre-training, to learn diverse human attributes from large-scale human images. Figure 2b shows the details. Compared to the human dance generation fine-tuning, the ControlNet branch for pose control is removed while the rest of the architecture remains the same. Consequently, we modify the objective as:
$L=\mathbb{E}_{\mathcal{E}(I),c_{f},\mu_{\theta}(g),\epsilon\sim\mathcal{N}(0,1),t}\left[\lVert\epsilon-\epsilon_{\theta}(z_{t},t,c_{f},\mu_{\theta}(g))\rVert_{2}^{2}\right].$
@@
Upon finishing, we initialize the U-Net and ControlNet branch for background control (highlighted with blue in Figure 2a) by the pre-trained model, and initialize the pose ControlNet branch with the pre-trained U-Net weight following [73], for human dance generation fine-tuning.
@@ 4.1 Experimental Setup
We train the models on the public TikTok dataset [24] for referring human dance generation. TikTok dataset consists of about 350 dance videos (with video length of 10-15 seconds) capturing a single-person dance. For each video, we first extract frames with 30fps, and run Grounded-SAM [30] and OpenPose [5] on each frame to infer the human subject mask for separating the foreground from the background and the pose skeleton. 335 videos are sampled as the training split. To ensure videos from the same person (same identity with same/different appearance) are not present in both training and testing splits, we collect 10 TikTok-style videos depicting different people from the web, as the testing split. We train our model on 8 NVIDIA V100 GPUs for 70K steps with image size $256\times 256$ and learning rate $2e^{-4}$. During training, we sample the first frame of the video as the reference and all others at 30 fps as targets. Both reference and target images are randomly cropped at the same position along the height dimension with the aspect ratio of 1, before resized to $256\times 256$. For evaluation, we apply center cropping instead of random cropping.
For human attribute pre-training, we use a combination of multiple public datasets (TikTok  [^1] We only use the training split for pre-training to avoid potential data leak. [24], COCO [35], SHHQ [10], DeepFashion2 [13], LAION [52]). We first run Grounded-SAM [30] with the prompt of “person” to automatically generate the human foreground mask, and then filter out images without human. This results in over 700K images for pre-training. All pre-training experiments are conducted on 4x8 NVIDIA V100 GPUs for 25K steps with image size $256\times 256$ and learning rate $1e^{-3}$. We initialize the U-Net model with the pre-trained weights of Stable Diffusion Image Variations [32]. The ControlNet branches are initialized with the same weight as the U-Net model, except for the zero-convolution layers, following [73]. After human attribute pre-training, we initialize the U-Net and ControlNet branch for background control by the pre-trained model, and initialize the pose ControlNet branch with the pre-trained U-Net weight, for human dance generation fine-tuning.
Figure 3: Visualizations of 5 representative scenarios for human image editing (best viewed when zoomed-in).
@@ 4.2 DisCo Applications
Benefiting from the strong human synthesis capability powered by the disentangled control design as well as human attribute pre-training, our DisCo provides flexible and fine-grained controlability and generalizability to arbitrary combination of human subject, pose and background. Given three existing images, each with distinct human subject, background and pose, there can be a total of 27 combinations. Here, we showcase 5 representatives scenarios in Figure 3 for human image editing: ($i$) Human Subject / Pose Re-targeting: the model have been exposed to training instances of the human subject or the ones of the pose, but the specific combinations of both are new to the model; ($ii$) Unseen Pose Generation: the human subject is from the training set, but the pose is novel, from the testing set; ($iii$) Unseen Human Subject Generation: the poses are sampled from the training set, but the human subject is novel, which is either from the testing set or crawled from the web; ($iv$) Unseen Pose & Human Subject Generation: both the human subject and pose are not present in the training set; ($v$) Full Unseen Composition: sampling a novel background from another unseen image/video based on the “unseen pose & human subject generation”. Examples in Figure 3 demonstrate that DisCo can flexibly update one of (or a composition of) human subject/background/pose in a given image to a user-specified one (or composition), either from existing training samples or novel images.
@@
Though it is not the focus of this paper, our final DisCo model can be readily and flexibly integrated with efficient fine-tuning techniques [23, 71, 26] for subject-specific fine-tuning on one or multiple images of the same human subject. We leave discussions and results of this setting to Appendix.
@@ 4.3 Main Results
We provide quantitative and qualitative comparisons against DreamPose [26], an image-to-video model designed for fashion domain with densepose control, which is the most relevant work to ours. DreamPose replaces the CLIP text feature with the image embedding with a dual CLIP-VAE encoder and adapter module. Instead of adopting ControlNet, it utilizes a sequence of denseposes [15] as the pose condition. It is worth noting that the TikTok dancing videos we evaluate on are all real-world user-generated content, which are much more complicated than those in existing dancing video datasets [67, 38, 34, 6, 26, 25], with clean background, same/similar clothing or fixed camera angles.
Quantitative Comparison. Since DisCo is applicable to both image and video generation for human dance synthesis, here we compare the models on both image- and video-wise generative metrics. To evaluate the image generation quality, we report frame-wise FID [17], SSIM [69], LISPIS [74], PSNR [22] and L1; while for videos, we concatenate every consecutive 16 frames to form a sample, to report FID-VID [1] and FVD [63]. As shown in Table 1, DisCo without human attribute pre-training (HAP) already significantly outperforms DreamPose by large margins across all metrics. Adding HAP further improves DisCo, reducing FID to $\sim 38$ and FVD to $\sim 440$. Not surprisingly, classifier-free guidance gives additional advantages to the generation quality of DisCo. The substantial performance gain against the recent SOTA model DreamPose evidently demonstrates the superiority of DisCo. Furthermore, we additionally collect 250 TikTok-style short videos from the web to enlarge the training split to $\sim$600 videos in total. The performance gain has shown the potential of DisCo to be further scaled-up.
Table 1: Quantitative comparisons of DisCo with the recent SOTA method DreamPose [26]. “CFG” and “HAP” denote classifier-free guidance and human attribute pre-training, respectively. $\downarrow$ indicates the lower the better, and vice versa. For DisCo ${}^{\dagger}$, we further scale up the fine-tuning stage to $\sim$600 TikTok-style videos.
Method Image Video
DreamPose [26] 79.46 0.509 28.04 0.450 6.91E-04 73.42 781.88
DreamPose [26] (CFG) 72.62 0.511 28.11 0.442 6.88E-04 53.36 671.50
DisCo (w/o HAP) 61.06 0.631 28.78 0.317 4.46E-04 32.56 555.44
DisCo (w/. HAP) 38.19 0.663 29.33 0.291 3.69E-04 20.24 439.94
DisCo (w/. HAP, CFG) 30.75 0.668 29.03 0.292 3.78E-04 18.86 393.34
DisCo ${}^{\dagger}$ (w/. HAP, CFG) 28.31 0.674 29.15 0.285 3.69E-04 15.77 348.04
@@
Figure 4: Qualitative comparison between our DisCo (w/ or w/o HAP) and DreamPose [26] on referring human dance video generation with the input of a reference image and a sequence of target poses. Note that the reference image and target poses are from the testing split, where the human subjects, backgrounds, poses are not available during the model training. Best viewed when zoomed-in.
4.4 Ablation Study
Architecture Design. Table 2 quantitatively analyzes the impact of different architecture
designs in DisCo. First, to ablate the control with reference image, we observe that either ControlNet or the cross-attention module struggle haodto handle the control of the whole reference image without disentangling the foreground from the background, leads to inferior quantitative results on most metrics. Though ControlNet-only baseline achieves better FVD scores, our visualizations in Figure 5 show that such architecture still struggles to maintain the consistency of human attributes and the stability of the background. For the encoding of reference foreground, DisCo with CLIP local feature produces better results than the one with CLIP global feature on 6 out of 7 metrics. We also explore to complement the CLIP local feature with VAE feature with a learnable adaptor following DreamPose [26], which leads to a slight better FID score but worse performance on all other metrics.
Table 2: Ablation on architecture designs without HAP. “ControlNet (fg+bg)” and “Attention (fg+bg)” in the first block denote inserting the control condition of reference image (containing both foreground and background) via a single ControlNet or cross-attention modules. “CLIP Global/Local” means using the global or local CLIP feature to represent the reference foreground. “CLIP Local + VAE” combines VAE features with CLIP Local features. Additional ablation results are included in Appendix.
@@ 4.4 Ablation Study
CLIP Global 63.92 0.621 28.61 0.311 5.00E-04 37.98 562.75
CLIP Local + VAE 59.74 0.623 28.52 0.331 4.79E-04 35.06 566.57
@@
N/A 0 61.06 0.631 28.78 0.317 4.46E-04 32.56 555.44
TikTok [24] 90K 50.68 0.648 28.81 0.309 4.27E-04 30.99 569.45
 + COCO [35] 110K 48.89 0.654 28.97 0.303 4.07E-04 26.68 516.89
 + SSHQ [10] 184K 44.13 0.655 29.00 0.300 3.93E-04 25.44 502.04
 + DpFashion2 [13] + LAION [52] 700K 38.19 0.663 29.33 0.291 3.69E-04 20.24 439.94
Pre-training Data Size. Table 3 investigates the effect of the data size in HAP stage by incrementally augmenting the pre-training data from open-source human image datasets. It is evident that a larger and more diverse pre-training data can yield better downstream results for referring human dance generation. Moreover, compared with “without pre-training” (1st row), adopting HAP on the same TikTok dataset as a self-supervised learning schema (2nd row) can already bring out significant performance gains. The final success of HAP comes from two-sides: 1) learning diverse human attributes from large-scale human image data; 2) the “easy-to-hard” training schema, with HAP focusing on reconstructing human images without pose editing, then learning pose control and implicit appearance distortions brought by motion in the fine-tuning stage.
@@ 5 Conclusion
Contributions. We revisit human dance synthesis in real-world for a more practical application and define a new problem setting: referring human dance generation with three key properties, faithfulness, generalizability, and compositionality. To tackle this problem, we propose DisCo, equipped with a novel architecture for disentangled control and an effective human attribute pre-training task. Extensive qualitative and quantitative results demonstrate the effectiveness of DisCo, which we believe is a step closer towards real-world applications for user-specific short video content generation.
Limitations. The limitation of DisCo sheds light on potential future directions: (1) incorporating hand keypoints for more fine-grained control of hand pose; (2) explicit temporal modeling to improve temporal consistency, especially when there are large motions; and (3) extending to more complicated scenarios, such as multi-person dance generation and human-object interaction.
Broader Impact. The proposed DisCo can be served as a strong starting point for real-world referring human dance generation, showing great potential for various applications, e.g., human image editing and video content creation. We believe that our proposed method will open up new possibilities for human-centric generation and editing, but also requires proper regulations to mitigate potential harm, such as the creation of deceptive content and infringement of human rights.
Appendix - DisCo: Disentangled Control for Referring Human Dance Generation in Real World

### v3

@@ 3.3 Human Attribute Pre-training
Table 1: Quantitative comparisons of DisCo with the recent SOTA method DreamPose. “CFG” and “HAP” denote classifier-free guidance and human attribute pre-training, respectively. $\downarrow$ indicates the lower the better, and vice versa. For DisCo ${}^{\dagger}$, we further scale up the fine-tuning stage to $\sim$600 TikTok-style videos. Methods with *** directly use target image as the input, including more information compared to the OpenPose.
Method Image Video
$*{}^{*}$ [52] 85.03 0.648 17.26 0.335 3.61E-04 90.09 405.22
$*{}^{*}$ [53] 54.47 0.672 18.14 0.296 3.21E-04 66.36 284.82
$*{}^{*}$ [73] 53.78 0.673 18.32 0.299 3.23E-04 72.55 306.17
DreamPose [25] 79.46 0.509 13.19 0.450 6.91E-04 80.51 551.56
DreamPose (CFG) 72.62 0.511 12.82 0.442 6.88E-04 78.77 551.02
DisCo (w/o HAP) 61.06 0.631 15.38 0.317 4.46E-04 73.29 366.39
DisCo (w/. HAP) 38.19 0.663 16.32 0.291 3.69E-04 61.88 286.91
DisCo (w/. HAP, CFG) 30.75 0.668 16.55 0.292 3.78E-04 59.90 292.80
DisCo ${}^{\dagger}$ (w/. HAP, CFG) 28.31 0.674 16.68 0.285 3.69E-04 55.17 267.75
Figure 3: Visualizations of 5 representative scenarios for human image editing. In the last column, we show that DisCo also generalizes to different image ratios and full-body human views. Check Figure 11 in Appendix for more results.
In utilizing the disentangled control architecture for DisCo, although it shows promises in pose control and background reconstruction, we find it remains challenging to have faithful generations with unseen human subject foregrounds and non-generic human views, demonstrating poor generalizability. The crux of this matter lies in the current training pipeline. It relies on high-quality human videos to provide training pairs of human images, with the same human foreground and background appearance, but different poses. Yet, we observe that current training datasets used for human dance generation confront a dilemma of “mutual exclusivity” — they cannot ensure both diversity in human attributes (such as identity, clothing, makeup, hairstyle, etc.) and the complicated poses due to the prohibitive costs of collecting and filtering human videos. As an alternative, human images, which are widely available over the internet, contain diverse human subject foregrounds and backgrounds, despite of the missing paired images with pose alterations.
This motivates us to propose a pre-training task, human attribute pre-training, to improve the generalizability and faithfulness. Rather than directly constructing the high-quality human dance video dataset, we explore an “easy-to-hard” training schema to step from image reconstruction to motion editing. Figure 0(b) shows the details. Specifically, we propose to reconstruct the whole image given the human foreground and background features. In this way, model effectively learns the distinguishment between human subject and foreground, as well as diverse human attributes from large-scale images. Compared to the human dance generation fine-tuning, the ControlNet branch for pose control is removed while the rest of the architecture remains the same. Consequently, we modify the objective as:
$L=\mathbb{E}_{\mathcal{E}(I),c_{f},\mu_{\theta}(g),\epsilon\sim\mathcal{N}(0,1),t}[\lVert\epsilon-\epsilon_{\theta}(z_{t},t,c_{f},\mu_{\theta}(g))\rVert_{2}^{2}].$ (3)
@@
For human dance generation fine-tuning, we initialize the U-Net and ControlNet branch for background control (highlighted with blue in Figure 0(a)) by the pre-trained model, and initialize the pose ControlNet branch with the pre-trained U-Net weight following [71].
@@ 4.1 Experimental Setup
We train the models on the public TikTok dataset [22] for referring human dance generation. TikTok dataset consists of about 350 dance videos (with video length of 10-15 seconds) capturing a single-person dance. For each video, we first extract frames with 30fps, and run Grounded-SAM [28, 34] and OpenPose [4] on each frame to infer the human subject mask and the pose skeleton. 335 videos are sampled as the training split. To ensure videos from the same person (same identity with same/different appearance) are not present in both training and testing splits, we collect 10 TikTok-style videos depicting different people from the web, as the testing split. We train our model on 8 NVIDIA V100 GPUs for 70K steps with image size $256\times 256$ and learning rate $2e^{-4}$. During training, we sample the first frame of the video as the reference and all others at 30 fps as targets. For equipping TM, we set $n=8$, and apply a learning rate of $5e^{-4}$ on the temporal convolution/attention layers. Both reference and target images are randomly cropped at the same position along the height dimension with the aspect ratio of 1, before resized to $256\times 256$. For evaluation, we apply center cropping instead of random cropping.
For human attribute pre-training, we use a combination of multiple public datasets (TikTok  [^1] We only use the training split for pre-training to avoid data leak. [22], COCO [32], SHHQ [9], DeepFashion2 [11], LAION [49]). We first run Grounded-SAM [28, 34] with the prompt of “person” to automatically generate the human foreground mask, and then filter out images without human. This results in over 700K images All pre-training experiments are conducted on 32 NVIDIA V100 GPUs for 25K steps with image size $256\times 256$ and learning rate $1e^{-3}$. We initialize the U-Net model with the pre-trained weights of SD Image Variations [24]. The ControlNet branches are initialized with the same weight as the U-Net model, except for the zero-convolution layers, following Zhang and Agrawala [71]. After human attribute pre-training, we initialize the U-Net and background ControlNet branch by the pre-trained model, and initialize the pose ControlNet branch with the pre-trained U-Net weight, for human dance generation fine-tuning.
@@ 4.2 DisCo Applications
Benefiting from the strong human synthesis capability powered by the disentangled control design as well as human attribute pre-training, our DisCo provides flexible and fine-grained controlability and generalizability to arbitrary combination of human subject, pose and background. Given three existing images, each with distinct human subject, background and pose, there can be a total of 27 combinations. Here, we showcase 5 representatives scenarios in Figure 3 for human image editing: ($i$) Human Subject / Pose Re-targeting: the model have been exposed to training instances of the human subject or the pose, but the specific combinations of both are new to the model; ($ii$) Unseen Pose Generation: the human subject is from the training set, but the pose is novel, from the testing set; ($iii$) Unseen Human Subject Generation: the poses are sampled from the training set, but the human subject is novel, which is either from the testing set or crawled from the web; ($iv$) Unseen Pose & Human Subject Generation: both the human subject and pose are not present in the training set; ($v$) Full Unseen Composition: further sampling a novel background from another unseen image/video based on $iv$ Examples in Figure 3 demonstrate that DisCo can flexibly update one of (or a composition of) human subject/background/pose in a given image to a user-specified one (or composition), either from existing training samples or novel images.
@@
Figure 4: Qualitative comparison between our DisCo (w/ or w/o HAP) and DreamPose on human dance video generation with the input of a reference image and a sequence of target poses. Note that the reference image and target poses are from the testing split, where the human subjects, backgrounds, poses are not available during the model training. Best viewed when zoomed-in.
Though it is not the focus of this paper, our final DisCo model can be readily and flexibly integrated with efficient fine-tuning techniques [21, 68, 25] for subject-specific fine-tuning on multiple images of the same human subject. We leave discussions and results of this setting to Appendix B.
@@ 4.3 Main Results
We provide quantitative and qualitative comparisons and test generalizability against both conventional motion transfer methods FOMM [52], MRAA [53], TPS [73], and most relevant work DreamPose [25], an image-to-video model designed for fashion domain with densepose control. It is worth noting that the TikTok dancing videos we evaluate on are all real-world content, which are much more complicated than those in existing dancing video datasets [63, 5, 23], with clean background and similar clothing.
@@
Quantitative Comparison. Since DisCo is applicable to both image and video generation for human dance synthesis, here we compare the models on both image- and video-wise generative metrics. To evaluate the image generation quality, we report frame-wise FID [15], SSIM [66], LISPIS [72], PSNR [20] and L1; while for videos, we concatenate every consecutive 16 frames to form a sample, to report FID-VID [1] and FVD [60]. As shown in Table 1, DisCo without human attribute pre-training (HAP) already significantly outperforms DreamPose by large margins across all metrics. Adding HAP further improves DisCo, reducing FID to $\sim 38$ and FVD to $\sim 280$.
Table 2: Video generation comparisons by adding the Temporal Modeling (TM). We employ HAP and CFG by default.
Method FID-VID $\downarrow$ FVD $\downarrow$
DisCo 59.90 292.80
DisCo (w/. TM) 34.19 260.34
DisCo ${}^{\dagger}$ 55.17 267.75
DisCo ${}^{\dagger}$ (w/. TM) 29.37 229.66
The substantial performance gain against the recent SOTA model DreamPose evidently demonstrates the superiority of DisCo. Furthermore, we additionally collect 250 TikTok-style short videos from the web to enlarge the training split to $\sim$600 videos in total. The performance gain has shown the potential of DisCo to be further scaled-up. As shown in Table 2, further incorporating temporal modeling to DisCo brings huge performance boosts to video synthesis metrics, e.g., improving FID-VID by $\sim 30$ and FVD by $\sim 60$.
Table 3: Ablation on architecture designs without HAP. “ControlNet (fg+bg)” and “Attention (fg+bg)” in the first block denote inserting the control condition of reference image (containing both foreground and background) via a single ControlNet or cross-attention modules. “CLIP Global/Local” means using the global or local CLIP feature to represent the reference foreground. “CLIP Local + VAE” combines VAE features with CLIP Local features. Additional ablation results are included in Appendix C.2.
@@
CLIP Global 63.92 0.621 14.58 0.311 5.00E-04 73.33 391.41
CLIP Local + VAE 59.74 0.623 14.84 0.331 4.79E-04 77.86 406.16
@@
N/A 0 61.06 0.631 15.38 0.317 4.46E-04 73.29 366.39
TikTok [22] 90K 50.68 0.648 15.62 0.309 4.27E-04 69.68 353.35
 + COCO [32] 110K 48.89 0.654 15.85 0.303 4.07E-04 62.15 326.88
 + SSHQ [9] 184K 44.13 0.655 16.09 0.300 3.93E-04 64.47 325.40
 + DpFashion2 [11] + LAION [49] 700K 38.19 0.663 16.32 0.291 3.69E-04 61.88 286.91
Beyond standard metrics, a user study with 50 unique participants was conducted. Participants rated images and videos 00 (worst) to $5$ (best). Figure 5 illustrates that DisCo (green histogram) clearly received higher user preference scores, with over 80% of users assigning a rating of $4$ or above. Further details on the study methodology and additional results are provided in the Appendix C.4.
Figure 5: Results of User Study: the overall quality score distribution on both synthesis (a) images and (b) videos of DreamPose and DisCo.
Generalizability. In addition to the TikTok data, we show strong zero-shot generation results of DisCo on several unseen out-of-distribution (OOD) data: Everybody Dance Now [5], AIST++ [31] and even wild youtube video in Figure 6.
Figure 6: Zero-shot synthesis of DisCo on several OOD data.
Furthermore, we quantitatively evaluate DisCo on two representative datasets in Table 5: conventional benchmark (TaiChi [52]) and OOD action recognition data (NTU-120 [33]). DisCo consistently outperforms the baselines.
4.4 Ablation Study
Architecture Design. Table 3 quantitatively analyzes the impact of different architecture designs in DisCo. First, to ablate the control with the reference image, we observe that either ControlNet or the cross-attention module struggles to handle the control of the whole reference image without disentangling the foreground from the background, leads to inferior quantitative results on most metrics. Though ControlNet-only baseline achieves better FVD scores, our visualizations in Figure 7 show that such architecture still struggles to maintain the consistency of human attributes and the stability of the background. For the encoding of the reference foreground, DisCo with CLIP local feature produces better results than the one with CLIP global feature on 6 out of 7 metrics. We also explore to complement the CLIP local feature with VAE feature with a learnable adaptor following DreamPose [25], which leads to a slight better FID score but worse performance on all other metrics.
Pre-training Data Size. Table 4 investigates the effect of the data size in HAP stage by incrementally augmenting the pre-training data from open-source human image datasets. It is evident that a larger and more diverse pre-training data can yield better downstream results for referring human dance generation. Moreover, compared with “without pre-training” (1st row), adopting HAP on the same TikTok dataset as a self-supervised learning schema (2nd row) can already bring out significant performance gains.
Table 5: Quantitative results on OOD human motion dataset.
Method TaiChi [52] NTU-120 [33]
FID $\downarrow$ FID-VID $\downarrow$ FVD $\downarrow$ FID $\downarrow$ FID-VID $\downarrow$ FVD $\downarrow$
FOMM 24.43 24.50 374.26 80.29 40.34 1439.50
MRAA 17.32 21.55 312.57 97.07 58.19 1441.79
DisCo 15.89 10.45 299.51 68.53 26.21 458.92
@@ 5 Conclusion
We revisit human dance synthesis for the more practical social media scenario and emphasize two key properties, generalizability and compositionality. To tackle this problem, we propose DisCo, equipped with a novel architecture for disentangled control and an effective human attribute pre-training task. Extensive qualitative and quantitative results demonstrate the effectiveness of DisCo, which we believe is a step closer towards real-world applications. Limitations. Though DisCo has shown promising results and decent human dance generation quality, it cannot handle hand posture well without the fine-grained hand pose control. Moreover, it is hard to be applied to multi-person scenarios and human-object interactions. Acknowledgement. This research/project is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-RP-2021-022).
Appendix - DisCo: Disentangled Control for Realistic Human Dance Generation